OpenStudy (anonymous):

Medal For Help!!!! Let f(x)= (x^2+3x-4) and g(x)=(x+4). a. Find f*g and state the domain. b. Find f/g and state the domain. Pleassee help

OpenStudy (anonymous):

domain of any polynomial is "all real numbers" \(\mathbb{R}\) or \((-\infty, \infty)\)

OpenStudy (anonymous):

for \(f\times g\) just stick a times sign between \(f\) and \(g\)

OpenStudy (anonymous):

for \(\left(\frac{f}{g}\right)\) make a fraction bar, put \(f\) on top and \(g\) in the bottom

OpenStudy (anonymous):

for the domain of \(\frac{f}{g}\) set \(x+4=0\), solve for \(x\) in one step, and then say "all real numbers except \(x=-4\)"

OpenStudy (anonymous):

OK. so for a. will it be (x^2+3x-4)(x+4), then I foil?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Coolsector @lala2 help pleassee